# Learning Multi-Modal Whole-Body Control for Real-World Humanoid Robots

### v1

@@ 4 Masked Humanoid Controller
5 Experimental Results
In this section, we present our evaluation protocol and simulation experiments to assess the MHC’s performance in executing diverse behaviors from full and partially-specified target motions. We compare the MHC with baseline approaches and analyze the impact of training dataset diversity as well as learning curriculum. Additionally, we discuss qualitative observations and the successful sim-to-real transfer of the MHC through real-world trials on the Digit V3 humanoid robot.
Motion Directive Dataset. We used IK retargetting to generate 75 kinematic motion trajectories selected from our motion sources as described in Section 4. We aimed to select a diverse set of motions, while avoiding aggressive motions that involve highly dynamic behavior such as jumping, kicking, and large sudden swinging of limbs. In order to test models on motions outside of their training sets, we divided the data into three sets, each containing a mix of motions from the original motion sources.
@@ 5 Experimental Results
While each set contains some basic motions related to hand waving and simple upper body movements, the general trend is an increasing level of difficulty as judged by the authors. We note that setC has a number of the most difficult motions, involving highly dynamic motion, such as tennis smashes and difficult boxing moves. The feasibility of these more difficult motions is unclear for our realistic Digit model.
Experiment Setup and Metrics. Each experiment involves evaluating an MHC controller on one of the retargeted motion datasets. For each motion we create 3 motion directives corresponding to a whole body directive, upper body command with lower body standing directive, upper body command with lower body locomotion directive. For each motion directive we evaluate the MHC over 200 episodes, each having a maximum time of 50 seconds (2500 steps). Each episode starts from a randomized point in the directive and then repeatedly cycles through the directive three times before resetting to another randomly selected starting point. Based on these episode runs we evauate the following metrics, which depend on the type of directive.
  • Failure Rate (Fail %): Percentage of episodes that result in failure (i.e. falling) before 50 seconds.
  • Mean per joint positional error ($\mathbb{E}_{\text{MPJPE}}$): Mean positional error for end effectors (hands, elbows, knees, feet) is calculated as the L2 norm of the current Euler position relative to the torso, compared to values from the directive. For partial directives, the error is computed only for unmasked joints. This is done for both partial and fully-specified directives before episode termination.
  • Root Drift (RootΔ): The mean of drift in root position during an episode from its commanded position when following partially or fully-specified directives. Computed as an L2 norm of current root position against expected root position in previous directive step.
Model Fail % $\mathbb{E}_{\text{MPJPE}}$ RootΔ
Locomotion+Offsets 49.213 0.156 0.168
LocomotionTrain 4.800 0.227 0.203
MHC 0.450 0.088 0.134
@@
Comparison to Baselines. There are no existing approach that handle the masked directive inputs of the MHC. Thus, we have developed two baselines to compare against: 1) Locomotion+Offset. This controller is trained on the first two curriculum stages and only sees the locomotion commands at its input. At inference time, it receives locomotion commands from the directives and for other unmasked joints directly uses the IK offsets in the target motions as the PD setpoints, 2) LocomotionTrain. This is the same as Locomotion+Offset during training and inference, but includes the IK offsets as part of the controller input. The motivation for LocomotionTrain is that by seeing the IK offsets that will be used during motion generation, the controller can adjust to account for the upcoming motion and be more robust.
We train the MHC and two baselines on the SetA data. The evaluation results are provided in Table 1. We see that the MHC outperforms the baselines across all metrics. Most significantly it achieves a failure rate of less than 1% which is significantly lower than the baselines. Interestingly, we see that LocomotionTrain achieves a significantly lower rate than Locomotion+Offsets, showing the utility having the controller learn to prepare for upcoming target motions.
@@
Table 2: Comparison of Models across Different Datasets. Note that higher failure rates tend to be biased toward smaller (better) values of the other metrics, since those metrics are only averaged over executions before failure.
Figure 2: Comparison of curriculum
stage 1 vs. training from scratch
Figure 3: Comparison of MHC vs. training whole-body directives without stability stage
Overall, these results show that our training approach can generalize to motions outside of the training set. However, more investigation is needed to understand the significance of training and testing with more difficult motions. In particular, the exploration of much longer training runs and continually monitoring of individual motion performance is likely to provide useful insights.
Curriculum Results. To evaluate the effectiveness of our curriculum design, we conducted an ablation study comparing the MHC trained with our complete curriculum against two variants: (1) NoCurriculum, a controller trained without any curriculum, and (2) LocoMask, a controller trained without the stabilization stage. All controllers were trained on setA. The results, shown in Figure 2 and Figure 3, demonstrate the importance of our structured curriculum. The NoCurriculum controller took significantly longer to learn and did not reach the same performance level as the MHC in the initial locomotion stage (Figure 2). The LocoMask controller exhibited greater instability, resulting in shorter average episode lengths (Figure 3). These findings highlight that our staged curriculum, which progressively learns locomotion, stability, and whole-body motion tracking, is very important for achieving our observed robustness and stability.
Sim-to-Real Results. We conduct a real-world demonstration of the MHC using the Digit V3 humanoid robot. Our supplementary video illustrates a number of successful examples and some failures. Some examples of successes include tracking of fully specified boxing directives, partial directives such as robust walking using joystick commands, box pick and place with locomotion + upper body partial directives as well as non-traditional zombie walking gaits. Failure modes included dancing motions that involved balancing on one feet for a substantial amount of time and fully extended arm poses for box pick up motions. Here holding the box outwards, away from the torso caused imbalance leading to a continuous drift from standing posture in order to stabilize. Wheile there is significant room to improvement these sim-to-real results demonstrate that the MHC framework is a promising path toward highly robust and versatile whole-body humanoid control.
@@
Figure 4: A) Locomotion Directives to follow joystick commands; B) Fully-specified directive for a boxing jab displaying whole body torso motion and feet coordination; C) Masked partial directives for upper body motion tracking and lower body locomotion commands to achieve the task of moving to a location, picking up the box, moving while holding the box and placing it at a location.
6 Limitations
One of the major current limitations is a significant sim-to-real gap. This is particularly prevalent for motions where the feet are placed widely or where one foot remains in the air for an extended period. Our hypothesis is that this sim-to-real limitation requires research that is more generally focused on sim-to-real transfer, possibly via the incorporation of real-world data. The MHC has been a useful vehicle for highlighting these gaps. Another limitation of our current training approach is that we only consider a limited set of masking patterns during training. Supporting fully general masking may lead to additional robustness and open up new possibilities for authoring behaviors. Finally, the MHC is currently does not account for interaction with external objects, such as the box, during a box pickup. Explicitly incorporating objects to be manipulated into the MHC input and training is an important direction for supporting practical applications.
Appendix A Dynamics Randomization
All models in our work were trained with a fixed dynamics randomization setup in order to transfer from simulation to real without learning the exact dynamics of the system. The ranges for all elements are given in Table 3.
Element Range(%)
Joint Damping [-50, 250]
Body Mass [-25, 25]
CoM Position [-5, 5]
Spring Stiffness [-10, 10]
Friction [-20, 20]
Range(radians)
Encoder Noise [-0.05, 0.05]
Ground Slope [-0.05, 0.05]
Table 3: Dynamics Randomization Ranges.
Appendix B Implementation Details
MHC has been trained using actor critic based PPO algorithm. Table 4 specifies the hyperparameters used for each training cycle independent of the curriculum.
Hyperparameter Value
Learning Rate (Actor & Critic) 3e-4
Buffer Size 50000
Number of Epochs 5
Batch Size (of episodes) 8
PPO Log-Clip Range 0.2
Maximum Gradient Norm 0.05
Discount Factor (Gamma) 0.95
GAE Lambda 0.95
Table 4: PPO Hyperparameters
Appendix C Reward structure
Table 5 gives a detailed view of our complete reward function. Note that style rewards are only ignored when following a fully-specified directive denoted by $I_{i}==1$.
Following notation has been used:
  • $i$ indicates current time step
  • $I_{i}==1$ indicated if all joints are being mimicked and its a whole body motion directive;
  • $v_{xy}$ = base linear velocity;
  • $s$ = current robot parameter;
  • $\boldsymbol{b}$ = base acceleration;
  • $c_{feet}$ = default foot position xyz and orientation rpy;
  • $qd(\cdot)$ = quaternion distance function;
  • $n_{c}$ = number of feet in contact with ground;
  • $\mathds{1}_{td}$ = boolean variable indicating a foot touchdown in the current timestep;
Reward Term Definition Weight Active if $I_{i}==1$
if $v_{i}==0$ [Standing] else [Walking]
Task $x,y$ velocity $\exp\left(-5\cdot(v_{xy})\right)$ $\exp\left(-5\cdot(v_{xy}-v_{i})^{2}\right)$ 0.15
Tracking Reward $\exp\left(-1.5\cdot||(\theta-\hat{\theta}_{i})\cdot I_{i}||\right)$ 0.2
Task, Style Yaw orient $\exp\left(-300\cdot qd(\omega_{z},w_{i})\right)$ 0.1
Roll, pitch orient $\exp\left(-30\cdot qd\big{(}\omega_{xy},(b_{i}[xy])\big{)}\right)$ 0.2
Base height $\exp\left(-20\cdot\big{|}s_{z}-(b_{i}[z])\big{|}\right)$ 0.05
Style Feet orientation $\exp\left(-\sum|s_{\text{feet, rpy}}-c_{\text{feet, rpy}}|\right)$ 0.05
Feet position $\exp\left(-3\cdot|S_{\text{feet, xyz}}-c_{\text{feet, xyz}}|\right)$ 1 0.05
Feet airtime 1 $\sum_{f\in(l,r)}(t_{air,f}-0.4)*\mathds{1}_{td,f}$ 1.0
Feet contact 1 $\begin{cases}1&\text{if }n_{c,t^{*}}=1\text{ for any }t^{*}\in[t-0.2,t]\\ 0&\text{else}\end{cases}$ 0.1
Arm $\exp\left(-3\cdot\big{|}\big{|}(\theta[\hat{J}]\text{ where }\hat{J}\in\text{ arms})\big{|}\big{|}\right)$ 0.03
Reg Base acceleration $\exp\left(-0.01\cdot\sum|b_{xyz}|\right)$ 0.1
Action difference $\exp\left(-0.02\cdot\sum|a_{t}-a_{t-1}|\right)$ 0.02
Torque $\exp\left(-0.02\cdot\frac{1}{N}\sum|t_{\text{motor}}|/t_{\text{max}}\right)$ 0.02
Table 5: Reward Terms

### v4

@@ Introduction
Experimental Results
We present our evaluation protocol and simulation experiments to assess the MHC’s performance. We compare the MHC with baseline approaches, analyze the impact of varying training sets, and learning curriculum. Additionally, we present qualitative and quantitative results of sim-to-real transfer to the Digit V3 robot.
Component Range
x velocity ($v_{ix}$) [-0.5, 2.0]m/s
y velocity ($v_{iy}$) [-0.5, 0.5]m/s
turn rate ($w_{i}$) [-0.5, 0.5]radians/s
torso x orientation ($b_{ix}$) [-0.314, 0.314]radians
torso y orientation ($b_{iy}$) [-0.314, 0.314]radians
torso height ($b_{ih}$) [0.5, 1.0]m
Standing mask bit ($I_{is}$) {0, 1}
Table 1: Predefined ranges for components of locomotion directives. Velocities ($v_{ix},v_{iy}$) and turn rate ($w_{i}$) are set to zero when the standing mask bit ($I_{is}$) is set to 11, indicating a standing directive.
Motion Directive Dataset
We used IK retargetting to generate 75 diverse kinematic motion trajectories. We selected this set of motions based on considering diversity, fit to the robot morphology, and avoiding overly aggressive motions such as jumping, kicking, and high velocity swinging of limbs. In order to study the generalization capabilities of our approach, we test MHC models on motions outside of their training sets, dividing the data into three sets, each containing a mix of motions from the original motion sources.
@@
While each set contains simple upper-body movements (such as waving), the general trend is an increasing level of difficulty as judged by the authors. We note that setC has a number of the most difficult motions, involving highly dynamic motion, such as tennis smashes and difficult boxing moves requiring precise and dynamic footwork combined with upper body motions. The feasibility of these more difficult motions is not obvious for our realistic Digit model.
Experimental Setup and Metrics
We evaluate the MHC on different input modalities including partial/whole-body mimicry, walking, and standing with perturbations. We use the full re-targeted dataset for whole-body mimicry and mask out the lower body targets for partial-body mimicry. Each episode for full/partial mimicry experiments command the complete motion directive once.
For pure walking experiments the directives are sampled from a predefined range as specified in Table 1. For each walking experiment episodes last for 10s and all non-locomotion directive components are masked out. Finally, for standing perturbation testing we follow prior work (van Marum et al. 2024) and provide an input directive of standing and apply varying forces at the base of the torso for a window of length of 25 steps to measure the root position drift. Below we describe the metrics used throughout our evaluations.
  • Mean End-Effector Positional Error ($\mathbb{E}_{\text{MEPE}}$): Mean positional error for end effectors (hands, elbows, knees, feet) is calculated as the L2 norm of the current Euler position relative to the torso, compared to values from the directive. For partial directives, the error is computed only for unmasked joints.
  • Root Drift (RootΔ): The mean of drift in root position during an episode from its commanded position when following partially or fully-specified directives. Computed as an L2 norm of the current root position against expected root position in the previous directive step.
  • Failure Rate (Fail %): Percentage of episodes that result in failure (i.e. falling).
Comparison to Baselines
Since no prior approaches handle masked directive inputs like the MHC, we compare our model with different baselines for locomotion and mimicry. For locomotion, we compare against a state-of-the-art standing and walking controller(van Marum et al. 2024). We also develop two natural baselines to compare against for locomotion + upper-body mimicry: 1)OffsetBlind. This model focuses on solely tracking locomotion components of input directives and does not receive upper-body IK offsets as model input. Rather it feeds upper-body offsets from the directive straight to the PD controllers without considering them in its locomotion planning. 2) OffsetAware. This model is the same as OffsetBlind, except that it receives upper-body offsets in its input. This allows it to adjust its locomotion control based on the anticipated upper body motions.
Model Walk RootΔ (m) Stand RootΔ (m)
1x m/s 0.5y m/s 50N 100N
MHC 0.413 0.343 0.256 0.311
SaW(SOTA) 0.803 0.388 0.109 0.452
Table 2: Locomotion performance comparison of MHC models with the SOTA Stand-and-Walk (SaW) controller. Root drift ($\mathrm{Root_{\Delta}}$) is measured for 10 s walking episodes and standing under applied torso forces.
Model Fail % $\mathbb{E}_{\text{MEPE}}$(in m) RootΔ(in m)
OffsetBlind 49.213 0.156 0.168
OffsetAware 4.800 0.227 0.203
MHC 0.450 0.098 0.141
Real World 5.0 0.117 -
@@
For this comparison experiment we train the MHC and all baselines on datset SetA. The results of our standing and walking evaluation in comparsions to the SOTA (van Marum et al. 2024) are provided in Table 2. We give results for the root drift when given walking commands (+1m/s x and +0.5m/s y) and for standing under different force perturbations (50N and 100N). The MHC demonstrates competitive performance in both walking and standing tasks, highlighting that MHC is on par with SOTA walking and standing capabilities.
Table 3 gives results comparing the MHC to the mimicry baselines. The MHC outperforms the baselines across all metrics. Most significantly it achieves a failure rate of less than 1% which is significantly lower than the baselines. Interestingly, we see that OffsetAware achieves a significantly lower failure rate than OffsetBlind, showing the utility of explicitly training the the controller to prepare for upcoming target motions provided in its input.
Tables 2 and 3 evaluate complementary aspects of MHC performance. Table 2 isolates locomotion robustness, comparing MHC against the state-of-the-art SaW controller under force perturbations. Table 3 focuses on whole-body imitation with locomotion integration, comparing MHC to mimicry baselines.
Evaluation of Generalization Performance
@@
Metric $\mathbb{E}_{\text{MEPE}}$(in m) RootΔ(in m)
setA setB setC setA setB setC
[…]
Table 4: Performance Comparison of MHC models on partial and whole-body directives computed on different datasets (setA, setB, setC)
Curriculum Results
We also train and evaluate $\text{MHC}_{\text{NC}}$ that is trained without any curriculum on the full set of directives (locomotion, partial and whole-body). We trained $\text{MHC}_{\text{NC}}$ until the training loss plateaued for a long period of time, providing substantially more training time than the MHC. In Table 5 we compare $\text{MHC}_{\text{NC}}$’s performance against MHC in perturbed standing, walking, and mimicry. We see that MHC clearly outperforms $\text{MHC}_{\text{NC}}$ on the tasks of walking and mimicry, which shows the effectiveness of our curriculum design. It is interesting to note that $\text{MHC}_{\text{NC}}$ slightly outperforms the MHC with respect to standing against perturbation. This appears to come with the trade-off of not aligning itself to the command directive as well as the MHC, particularly in the lower body.
@@
Figure 2: Real-world demonstrations of our approach. A) Locomotion Directives using joystick commands. B) Fully-specified directive for a boxing jab; showing whole-body torso motion and foot coordination. C) Handcrafted sequence of masked directives combining upper body motion trajectory and lower body joystick commands; showing the ability to move to a location, pick up a box, move while holding it, and place it at a target location.
Metric MHC $\text{MHC}_{\text{NC}}$
Stand (50 N) RootΔ(in m) 0.256 0.189
(100 N) RootΔ(in m) 0.311 0.234
Walk (1x m/s) RootΔ(in m) 0.413 1.346
(0.5y m/s) RootΔ(in m) 0.343 0.937
$\mathbb{E}_{\text{MEPE}}$(in m) 0.098 0.149
Mimic RootΔ(in m) 0.141 0.327
Fail % 0.450 4.875
Table 5: Performance metrics comparing a model trained with curriculum (MHC) against a model trained without curriculum ($\text{MHC}_{\text{NC}}$) on the tasks of standing, walking and mimicry (averaged for both partial and whole-body directives from motions in dataset setA).
Sim-to-Real Results
We conducted real-world demonstrations of the MHC using the Digit V3 humanoid robot. Figure 2 shows trials including box locomanipulation, which involved the input modality of locomotion and upper-body end-effector tracking. For a quantitative evaluation we tested the 20 motions in setA on the real robot, performing 3 repetitions of each motion. The results are given in Table 3. While the failure rate (5%) and joint position error are higher than observed in simulation, these results indicate our approach is able to achieve highly non-trivial sim-to-real transfer.
Future Work
This work motivates a number of important research directions: 1) Addressing the remaining sim-to-real gaps, especially for motions with wide foot placements or extended single-foot balancing stances. Incorporating real-world data into the sim-to-real process is likely to play an important role. 2) Continuing to increase the number and types of motions used for MHC training. Currently, for Digit V3, the bottleneck is our reliance on the MuJoCo simulator, compared to other higher throughput simulators such as Isaac Sim, which do not yet support the Digit V3 model due to modeling complexities. 3) Increasing the variety of masking patterns used for training to further enhance versatility. 4) Handling inconsistent and/or impossible input directives (e.g. root velocity is inconsistent with other joint components), which are currently not part of training.
Acknowledgments
This work is supported by NSF Award 2321851 and DARPA contract HR0011-24-9-0423.